the area of a rectangular garden is 42 square meters. the width is 5 meters longer than one-third of the length. find the length and the width of the garden
Answer by gsmani_iyer(201)


Let the length of the rectangle = x
So, the breadth of the rectangle = (x/3)+5
Area = length * breadth.=> 42 Sq.mtr.
= x * [(x/3) + 5] = 42 Sq.mtrs.
= x^2/3 + 5x = 42 Sq.mtrs.
= x^2/3 + 5x - 42 = 0 ......... ( 1 ). Now Eqn.1 * 3, we get
= x^2 + 15x - 126 = 0 ......... ( 2 )
= (x + 21)(x - 6) = 0 So either (x+21) or (x-6) = 0
If x+21 = 0, then x=-21 whereas the length cannot be negative. So,
x = 6. [(x/3)+5] = 7.
Length of the rectangle = 6 mtrs.; breadth = 7 mtr.;
I hope this is clear to you. All the best. Feel free to contact me if you have doubts in maths.
gsm